OpenStudy (anonymous):

Anna is buying 2/4 of a pound of beans for dinner. Mercy is buying 7/4 the amount of beans that Anna is buying. How many pounds of beans is Mercy buying?

OpenStudy (anonymous):

Should be this way: Anna buys= 2/4 or 1/2 of a pound of beans. Mercy buys=7/4(amount of beans Anna buys) Which is =1/2(7/4) = 7/8

OpenStudy (anonymous):

ok thnx :) bro